CHW-10 ch-7 Problem 7.25 Review Part A How much work was done on the pinecone by air resistance? Express your answer using two significant figures A 0.16 kg pinecone falls 17 m to the ground, where it lands with a speed of 13 m/s. W15 Submit X Incorrect; Try Again; 14 attempts remaining Part B What was the average force of air resistance exerted on the pinecone? Express your answer using two significant figures Submit Request Answer
Answer #1

here,

mass of pinecone , m = 0.16 kg

height , h = 17 m

the final speed , v = 13 m/s

a)

the work done by the air resistance on pinecone , W = change in total energy

W = m * g * h - 0.5 * m *v^2

W = 0.16 * 9.81 * 17 - 0.5 * 0.16 * 13^2

W = 13.16 J

b)

let the average air resistance force be F

work done , W = F * h

13.16 = F * 17

solving for F

F = 0.77 N

the average air resistance force is 0.77 N